OpenStudy (anonymous):

Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@FEARLESS_RAIN medal and fan?

OpenStudy (anonymous):

How did you get that, can you explain please?

OpenStudy (anonymous):

The first one because 2(4) - 5(1) = 3 and (4) - 3(1) = 1 Just plug in the numbers where you see x and y

OpenStudy (anonymous):

8 - 5 = 3 and 4 - 3 = 1 ?

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

i need to refresh my page

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(x,y) = (4, 1) so whenever you see an x you replace it with 4 and with y you replace it with 1. Then multiply to get an answer

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2x - 5y = 3 }\) \(\large\color{black}{ \displaystyle x - 3y = 1 }\) (step 1) multiply the second equation times -2. (step 2) THEN add the equations.

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

alrighty, what do you get ?

OpenStudy (solomonzelman):

what do you get for step 1 ?